OpenStudy (anonymous):

Given the function f(x)=(x-2)/(x-5), determine an interval and a point when the average rate of change and instantaneous rate of change are equal

OpenStudy (anonymous):

why? :)

OpenStudy (anonymous):

ooh i can't read. determine an interval AND a point sorry

OpenStudy (anonymous):

this is an open ended question. pick an interval, one that does not include 5 because the function is not defined there. pick any interval, but make it easy

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

so you want to pick or do you want me to pick?

OpenStudy (anonymous):

u could pick it :) doesn't matter

OpenStudy (anonymous):

it really doesn't matter, i would pick something simple, like say \([6,8]\)

OpenStudy (anonymous):

we have some computation to do, namely \[f(8)=\frac{8-2}{8-5}=\frac{6}{3}=2\] and \[f(6)=\frac{6-2}{6-5}=\frac{4}{1}=4\] see, nice and easy

OpenStudy (anonymous):

now we compute \[\frac{f(8)-f(6)}{8-6}=\frac{2-4}{2}=-1\]

OpenStudy (anonymous):

that is the average rate on that interval. last job is to take the derivative, set it equal to \(-1\) and solve for \(x\) that is where the average rate of change will be equal to the instantaneous rate

OpenStudy (anonymous):

but what is the derivative ? sorry i don't know it ?

OpenStudy (anonymous):

derivative is \[f'(x)=-\frac{3}{(x-5)^2}\] so your final job is to set \[-\frac{3}{(x-5)^2}=-1\] and solve for \(x\) should be ok now right?

OpenStudy (anonymous):

wait hold the phone. if you do not know the derivative, then you cannot know the instantaneous rate of change. can't know what it means

OpenStudy (anonymous):

i know the instantaneous rate of change :)

OpenStudy (anonymous):

ok so you know the derivative, yes?

OpenStudy (anonymous):

your final job in this one is to solve \[-\frac{3}{(x-5)^2}=-1\] \[-3=-(x-5)^2\] \[(x-5)^2=3\] \[x-5=\pm\sqrt{3}\] \[x=5\pm\sqrt{3}\] the one that is in the interval \([6,8]\) is \(5+\sqrt{3}\)

OpenStudy (anonymous):

wait ?

OpenStudy (anonymous):

Do you know what a "derivative" is?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

never mind, i understand it :) thanks a lottttttttt (satellite73):)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Nice job, @satellite73 , I don't know why you don't have a 100.

OpenStudy (anonymous):

my score decrease, probably from every time i make a typo

OpenStudy (anonymous):

hihi, u alway help me out :)

OpenStudy (anonymous):

actually i don't know why it was 100 and it went down. frankly i would perfer to remain undgraded

OpenStudy (anonymous):

glad to help

OpenStudy (anonymous):

Well, forgive me if I'm not that bothered by your 99. :P It's still an A+.